Given two integers a and b, your task is to calculate and print the following four values:-

a+b

a-b

a*b

a/b using python


1
Expert's answer
2022-05-16T16:07:52-0400


print("------------------ My Little calculator ----------------")
a=int(input('a='))
b=int(input('b='))
print("a+b="+str(a+b))
print("a-b="+str(a-b))
print("a*b="+str(a*b))
print("a/b="+str(a/b))